### v1

@@ Abstract
Neutrinos are elementary particles that interact only very weakly with matter. Neutrino experiments are therefore usually big, with masses on the multi-ton scale. The thresholdless interaction of coherent elastic scattering of neutrinos on atomic nuclei leads to drastically enhanced interaction rates, that allows for much smaller detectors. The study of this process gives insights into physics beyond the Standard Model of particle physics. The Conus+ experiment was designed to first detect elastic neutrino-nucleus scattering in the fully coherent regime with low-energy neutrinos produced in nuclear reactors. For this purpose, semiconductor detectors based on high-purity germanium crystals with extremely low energy threshold of 160180 eV were developed. Here we show the first observation of a neutrino signal with a statistical significance of 3.7 sigma from the Conus+ experiment, operated at the nuclear power plant in Leibstadt, Switzerland. In 119 days of reactor operation (395106) neutrinos were measured compared to a predicted number from simulations assuming standard model physics of (34734) events. The good agreement between data and prediction constrains many parameters in various theoretical models. With increased precision, there is potential for fundamental discoveries in the future. The Conus+ results in combination with other measurements of this interaction channel might therefore mark a starting point for a new era in neutrino physics.
@@ Introduction
Neutrino sources suitable for the measurement of the coherent elastic neutrino-nucleus scattering (CEvNS) need to have energies of a few up to tens of MeV. Intense neutrino sources fulfilling this condition are, for example, the Sun, supernova explosions, accelerator-based sources, or nuclear reactors. It took 43 years after its prediction until CEvNS was first detected in 2017 by the COHERENT experiment in a scintillating crystal of cesium iodide [2]. Here, the neutrinos are generated at a Spallation Neutron Source (SNS) when pions decay at rest. At the same neutrino source the CEvNS measurement was confirmed with argon [3] and germanium [4] as target materials. Nuclear reactors emit neutrinos of lower energies than SNS sources. This has the advantage of a strongly enhanced sensitivity for several parameters in beyond the standard model (BSM) theories. However, it makes the detection of neutrinos much harder because of the requirement of extremely low energy thresholds in the detector.
There is a long history of successful experiments using reactor antineutrinos as sources, including the first neutrino detection in 1956 [5]. An intense antineutrino flux is produced after the fission process during the beta decay of the fission fragments. Nuclear reactors can be considered as well-defined, pure and point-like sources of specific neutrino flavor type, the electron antineutrinos. In recent years, reactor experiments allowed to study neutrino oscillation parameters [6, 7, 8, 9] and constrain the existence of sterile neutrinos [10, 11, 12, 13]. Today, there is a worldwide effort to measure CEvNS close to nuclear reactors [14, 15, 16, 17, 18, 19, 20, 21, 22, 23]. In terms of neutrino oscillation studies, it is a complementary approach, since CEvNS is sensitive to all known neutrino flavors, the electron, muon and tau neutrinos/antineutrinos, whereas the inverse beta decay as standard detection technique for reactor neutrinos is only sensitive to electron antineutrinos. In addition, there is a wide range of studies and topics that can be addressed in CEvNS measurements. For example, they can be used to probe nuclear structures, they are highly relevant in astrophysics, and they also provide important input for present and future dark matter experiments, which are limited in sensitivity by the CEvNS rate of solar neutrinos. Very recently, first indications for such a CEvNS signal were found in dark matter experiments with xenon as target [24, 25]. Moreover, Conus+ technology has the potential to use neutrinos for reactor monitoring and safeguard applications in the future.
@@
The Conus experiment [18] began operating in 2018 at the nuclear power plant in Brokdorf, Germany (KBR). There, the experiment took data until 2022 at a distance of only 17.1 m from the center of the reactor core. At the maximum thermal power of the reactor of 3.9 GW, the corresponding flux at the detector position was estimated to be 1013 antineutrinos per cm2 and second. The experimental setup consisted of four high-purity germanium (HPGe) detectors. Each diode had a mass of about 1 kg leading to a total fiducial germanium mass of kg [26]. As a final result of the measurement at KBR, the neutrino flux was constrained with 90% confidence level to a factor 1.6 above the signal expectation [27]. This world-best upper limit on the CEvNS interaction rate at nuclear reactors so far allowed us to exclude significant deviations from the SM or to test the standard description of signal quenching due to dissipation effects in the germanium material in the energy region of interest.
@@
The search for CEvNS in nuclear reactors is a challenging task for various reasons. Since a high neutrino flux is required, positions close to the reactor core inside the reactor building are preferred. The environment inside this inner control zone is quite different from the working conditions in a common research laboratory. There are several restrictions related to the materials allowed, earthquake safety, access, data transfer, etc. Appropriate solutions on all these topics were found in close cooperation with the KBR and KKL staff. In addition, there is limited protection against cosmic radiation, as the overburden at the CONUS+ site corresponds only to the equivalent 7.4 m water. In general, the radioactivity level has to be kept under control to perform successful rare event searches. For example, cosmic muons produce electromagnetic cascades and neutrons in the building structure and shield materials. These cascades can create event signatures similar to neutrinos in the HPGe detectors. Mitigation of detector signals created by such cosmic radiation or environmental radioactivity (background events) is achieved by using an effective shield structure around the detectors [28] as depicted in Fig.1. Furthermore, the energy of the nuclear recoils after neutrino scattering is very low. The unit used for the energy measured by the Ge detectors is given in eV ( J) and should be interpreted as ionization energy. It was a longstanding effort to reach the required threshold levels in the detectors. Our currently lowest threshold level of 160 eV is only two orders of magnitude above the typical semiconductor band gap, which defines the minimum energy to create one electron-hole pair.
@@
The reactor correlated background events, in particular the ones from neutrons produced in the reactor, are thoroughly studied in [30] and found to be negligible. A significant background contribution spectrum shown in Fig.2 on the left is from neutrons, which are induced by cosmic rays in the materials of the Conus+ shield. In addition, there is a relevant fraction of direct neutrons up to 100 MeV energy already produced in the atmosphere in cosmic ray air showers. Other components contributing to the spectrum are metastable states in germanium and the radioactive noble gas radon. The background during reactor off was lower since a vessel lid with a thickness of a few cm of steel was positioned right above the Conus+ room during reactor outage providing additional overburden. According to simulations, this reduction was 3% for muons and 19% for direct cosmic neutrons. Variations in the radon level inside the detector chamber were corrected. Otherwise, no significant differences are observed during reactor on and off. The dominant background contributions, such as the muon-induced or neutron background, are identical for the three detectors used in the analysis.
The calculation of the signal prediction is based on a method proposed in [31]. The neutrino flux depends on the fission fractions of the uranium and plutonium isotopes in the reactor: 235U, 238U, 239Pu and 241Pu, respectively. The average contribution of these isotopes to the flux during data collection with reactor on is 53%, 8%, 32% and 7%. In principle, there is no energy threshold for the CEvNS interaction itself and therefore there is a potential to measure neutrinos even below the threshold for the inverse beta decay reaction of about 1.8 MeV. Beyond the reactor conditions and its thermal power, the neutrino rate measured in the detector strongly depends on dissipation processes in the germanium crystals. The ionization energy observed in the detector is reduced with respect to the deposited recoil energy, a characteristic known as quenching. There was a debate, if this quenching factor is enhanced at low energies as compared to the Lindhard theory [32, 33]. From Conus data, there is no indication of any deviation from the Lindhard model [34, 27].
@@
The energy threshold of the detectors is estimated for each of them individually [28]. The lowest value for the C3 detector is only 160 eV. For the other two detectors used in the analysis, the thresholds after energy non-linearity correction were set slightly above, at 170 eV and 180 eV. One of the four detectors (C4) showed significant instabilities in the rate above 250 eV and was therefore removed from the data set. The thresholds were defined in a way to assure that contributions from electronic noise and microphonics are negligible in the region of interest. The trigger efficiency was determined for each detector using a pulse generator and was found to be close to 100% above the thresholds.
Another crucial requirement for detecting CEvNS at a nuclear reactor is the stability of environmental parameters, electronic noise and background rates. For example, temperature fluctuations can induce cryocooler power variations, which might create microphonic events. Microphonic noise and rate correlations with room temperature were further reduced compared to previous analyses [18, 27] by an improved cooling system [28]. The stability of the detector parameters such as energy resolution or trigger efficiency was regularly checked with the pulse generator. Variations of the noise peak were carefully monitored and data were only selected for the analysis in case they were below a defined level.
The data set used in the analysis reported here includes reactor on periods between November 2023 and July 2024 (347 kg d) and a off period during reactor outage in May 2024 (63 kg d). These data sets are fitted simultaneously for the individual detectors and the combination of the three detectors in an energy window between 160 eV to 800 eV. The signal is extracted based on a profile likelihood ratio test. Systematic uncertainties are considered with Gaussian pull terms. The data acquisition system in principle allows for background rejection studying the shape of the digitized pulses [35]. This option was not yet applied in the current analysis, but it is planned to be used in future analyses.
| Detector | Energy threshold (eV) | Signal events data | Neutrinos predicted | Ratio |
| C2 | 180 | |||
| C3 | 160 | |||
| C5 | 170 | |||
| combined | ||||
[…]
From the combined fit, a neutrino signal of () events in the reactor on data set was obtained. The significance of this event excess corresponds to . The neutrino signal at low energies of the spectrum is illustrated in Fig.3. Fits using just single detectors independently give consistent results as shown in Table 1. The result was cross-checked in two implementations of the likelihood fit and agreed within less than 2%. Systematic uncertainties related to background model, non-linearity correction and fit systematics are included in the combined fit result.
@@
The Conus+ signal of () measured neutrinos is fully consistent with the expectation of () events. This implies agreement of the Conus+ data with the CEvNS cross-section of the SM and the estimated antineutrino flux based on the thermal power of the reactor. Moreover, the detected rate is in very good agreement with the predicted Ge quenching using the Lindhard theory with a quenching parameter as measured in [34]. The deviations from Lindhard theory claimed in [33] and the excess reported in [17], which is based on it, are both ruled out by this result. The claim was already disfavored before by the independent measurement of the quenching factor [34], previous CONUS results [27] and theoretical considerations [36].
@@
In Table 2, the ratio of measured and predicted neutrino interactions is shown for the Conus+ result. This ratio is compared to the ratios measured in the COHERENT measurements at higher neutrino energies using multiple target nuclei and the results from coherent scattering of solar neutrinos. The use of different target nuclei allows to study the quadratic enhancement of the cross-section by the number of neutrons in the target. In the COHERENT data for Ge, the measured rate was found slightly below the predicted value, although not highly significant. Such a deficit was not confirmed in the Conus+ data. The combination of the CEvNS result of Conus+ with results of the germanium target data of the COHERENT experiment can in principle be used to extract information about nuclear form factors in neutrino light.
After the first detection of CEvNS at a nuclear reactor, as reported in this article, the next step will be a more precise measurement of the CEvNS cross-section. Higher precision can be achieved by increasing the target mass, lowering the energy threshold of the detectors and longer operation times, in particular in the reactor off phases. Therefore, 3 of the 4 Conus+ detectors were replaced in November 2024 by a newer generation. These new detectors have a larger mass of 2.4 kg each. First characterizations in the MPIK laboratory indicated even lower energy thresholds. The detector with the lowest energy threshold in the first Conus+ run (C3) was kept as a reference for a better comparison between the phases of the experiment. With this detector configuration it is planned to measure for another few years.
A high-statistics CEvNS measurement might open a new phase in fundamental physics and will allow to study physics within and beyond the SM. The measured CEvNS rate can for example be affected by new mediator particles similar to the Z boson, electromagnetic properties of neutrinos, or non-standard interactions. Moreover, it is possible to study the Weinberg angle at low energies, the existence of sterile neutrinos or supernovae astrophysics. There are also connections to dark matter experiments. A precise CEvNS measurement will allow to learn more about the neutrino sources as the Sun or nuclear reactors. The evolution of the reactor thermal power and fissile isotope concentrations in fuel elements could be monitored with rather small and mobile neutrino detectors. In summary, there is a wide range of topics ranging from BSM theories, nuclear physics and astrophysics that can be addressed with CEvNS measurements at nuclear reactors. The Conus and Conus+ experiments are pioneering in this field.

### v2

@@ Abstract
Neutrinos are elementary particles that interact only very weakly with matter. Neutrino experiments are therefore usually big, with masses on the multi-ton scale. The thresholdless interaction of coherent elastic scattering of neutrinos on atomic nuclei leads to drastically enhanced interaction rates, that allows for much smaller detectors. The study of this process gives insights into physics beyond the Standard Model of particle physics. The Conus+ experiment was designed to first detect elastic neutrino-nucleus scattering in the fully coherent regime with low-energy neutrinos produced in nuclear reactors. For this purpose, semiconductor detectors based on high-purity germanium crystals with extremely low energy threshold of 160180 eV were developed. Here we show the first observation of a neutrino signal with a statistical significance of 3.7 sigma from the Conus+ experiment, operated at the nuclear power plant in Leibstadt, Switzerland. In 119 days of reactor operation (395106) neutrinos were measured compared to a predicted number from calculations assuming standard model physics of (34759) events. The good agreement between data and prediction constrains many parameters in various theoretical models. With increased precision, there is potential for fundamental discoveries in the future. The Conus+ results in combination with other measurements of this interaction channel might therefore mark a starting point for a new era in neutrino physics.
@@ Introduction
Neutrino sources suitable for the measurement of the coherent elastic neutrino-nucleus scattering (CENS) need to have energies of a few up to tens of MeV. Intense neutrino sources fulfilling this condition are, for example, the Sun, supernova explosions, accelerator-based sources, or nuclear reactors. It took 43 years after its prediction until CENS was first detected in 2017 by the COHERENT experiment in a scintillating crystal of cesium iodide [2]. Here, the neutrinos are generated at a Spallation Neutron Source (SNS) when pions decay at rest. At the same neutrino source the CENS measurement was confirmed with argon [3] and germanium [4] as target materials. Nuclear reactors emit neutrinos of lower energies than SNS sources. This has the advantage of a strongly enhanced sensitivity for several parameters in beyond the standard model (BSM) theories. However, it makes the detection of neutrinos much harder because of the requirement of extremely low energy thresholds in the detector.
There is a long history of successful experiments using reactor antineutrinos as sources, including the first neutrino detection in 1956 [5]. An intense antineutrino flux is produced after the fission process during the beta decay of the fission fragments. Nuclear reactors can be considered as well-defined, pure and point-like sources of specific neutrino flavor type, the electron antineutrinos. In recent years, reactor experiments allowed to study neutrino oscillation parameters [6, 7, 8, 9] and constrain the existence of sterile neutrinos [10, 11, 12, 13]. Today, there is a worldwide effort to measure CENS close to nuclear reactors [14, 15, 16, 17, 18, 19, 20, 21, 22, 23]. In terms of neutrino oscillation studies, it is a complementary approach, since CENS is sensitive to all known neutrino flavors, the electron, muon and tau neutrinos/antineutrinos, whereas the inverse beta decay as standard detection technique for reactor neutrinos is only sensitive to electron antineutrinos. In addition, there is a wide range of studies and topics that can be addressed in CENS measurements. For example, they can be used to probe nuclear structures, they are highly relevant in astrophysics, and they also provide important input for present and future dark matter experiments, which are limited in sensitivity by the CENS rate of solar neutrinos. Very recently, first indications for such a CENS signal were found in dark matter experiments with xenon as target [24, 25]. Moreover, Conus+ technology has the potential to use neutrinos for reactor monitoring and safeguard applications in the future.
@@
The Conus experiment [18] began operating in 2018 at the nuclear power plant in Brokdorf, Germany (KBR). There, the experiment took data until 2022 at a distance of only 17.1 m from the center of the reactor core. At the maximum thermal power of the reactor of 3.9 GW, the corresponding flux at the detector position was estimated to be 1013 antineutrinos per cm2 and second. The experimental setup consisted of four high-purity germanium (HPGe) detectors. Each diode had a mass of about 1 kg leading to a total fiducial germanium mass of kg [26]. As a final result of the measurement at KBR, the neutrino flux was constrained with 90% confidence level to a factor 1.6 above the signal expectation [27]. This world-best upper limit on the CENS interaction rate at nuclear reactors so far allowed us to exclude significant deviations from the SM or to test the standard description of signal quenching due to dissipation effects in the germanium material in the energy region of interest.
@@
The search for CENS in nuclear reactors is a challenging task for various reasons. Since a high neutrino flux is required, positions close to the reactor core inside the reactor building are preferred. The environment inside this inner control zone is quite different from the working conditions in a common research laboratory. There are several restrictions related to the materials allowed, earthquake safety, access, data transfer, etc. Appropriate solutions on all these topics were found in close cooperation with the KBR and KKL staff. In addition, there is limited protection against cosmic radiation, as the overburden at the CONUS+ site corresponds only to the equivalent 7.4 m water. In general, the radioactivity level has to be kept under control to perform successful rare event searches. For example, cosmic muons produce electromagnetic cascades and neutrons in the building structure and shield materials. These cascades can create event signatures similar to neutrinos in the HPGe detectors. Mitigation of detector signals created by such cosmic radiation or environmental radioactivity (background events) is achieved by using an effective shield structure around the detectors [28] as depicted in Fig.1. Furthermore, the energy of the nuclear recoils after neutrino scattering is very low. The unit used for the energy measured by the Ge detectors is given in eV ( J) and should be interpreted as ionization energy. It was a longstanding effort to reach the required threshold levels in the detectors. Our currently lowest threshold level of 160 eV is only two orders of magnitude above the typical semiconductor band gap, which defines the minimum energy to create one electron-hole pair.
@@
The dominant background contributions, such as the muon-induced or neutron background, are identical for the three detectors used in the analysis. A comparison between background model and data is shown in Fig.2. In addition to the neutrons, which are created by muons in the materials of the Conus+ shield, there is a relevant fraction of direct neutrons up to 100 MeV energy already produced in the atmosphere in cosmic ray air showers. The background during reactor off was lower since a vessel lid with a thickness of a few cm of steel was placed right above the Conus+ room during reactor outage providing additional overburden. According to simulations, this reduction was 3% for muons and 19% for direct cosmic neutrons. Other components contributing to the background spectrum are from metastable states in germanium and the radioactive noble gas radon. Variations in the radon level inside the detector chamber were corrected. Otherwise, no significant differences are observed during reactor on and off. The reactor correlated background events, in particular the ones from neutrons produced in the reactor, are thoroughly studied in [30] and found to be negligible.
The calculation of the signal prediction is based on a method proposed in [31]. The neutrino flux depends on the fission fractions of the uranium and plutonium isotopes in the reactor: 235U, 238U, 239Pu and 241Pu, respectively. The average contribution of these isotopes to the flux during data collection with reactor on is 53%, 8%, 32% and 7%. In principle, there is no energy threshold for the CENS interaction itself and therefore there is a potential to measure neutrinos even below the threshold for the inverse beta decay reaction of about 1.8 MeV. Beyond the reactor conditions and its thermal power, the neutrino rate measured in the detector strongly depends on dissipation processes in the germanium crystals. The ionization energy observed in the detector is reduced with respect to the deposited recoil energy, a characteristic known as quenching. There was a debate, if this quenching factor is enhanced at low energies as compared to the Lindhard theory [32, 33]. From Conus data, there is no indication of any deviation from the Lindhard model [34, 27].
@@
The energy threshold of the detectors is estimated for each of them individually [28]. The lowest value for the C3 detector is only 160 eV. For the other two detectors used in the analysis, the thresholds after energy non-linearity correction were set slightly above, at 170 eV and 180 eV. One of the four detectors (C4) showed significant instabilities in the rate and was therefore removed from the data set. The thresholds were defined in a way to assure that contributions from electronic noise and microphonics are negligible in the region of interest. The trigger efficiency was determined for each detector using a pulse generator and was found to be close to 100% above the thresholds.
Another crucial requirement for detecting CENS at a nuclear reactor is the stability of environmental parameters, electronic noise and background rates. For example, temperature fluctuations can induce cryocooler power variations, which might create microphonic events. Microphonic noise and rate correlations with room temperature were further reduced compared to previous analyses [18, 27] by an improved cooling system [28]. The stability of the detector parameters such as energy resolution or trigger efficiency was regularly checked with the pulse generator. Variations of the noise peak were carefully monitored and data were only selected for the analysis in case they were below a defined level.
The data set used in the analysis reported here includes reactor on periods between November 2023 and July 2024 (327 kg d) and a off period during reactor outage in May 2024 (60 kg d). The data sets of the three detectors are fitted simultaneously in an energy window between 160 eV and 800 eV. The signal is extracted based on a profile likelihood ratio test. Systematic uncertainties are considered with Gaussian pull terms. The data acquisition system in principle allows for background rejection by studying the shape of digitized pulses [35]. This option was not yet applied in the current analysis, but it is planned to be used in future analyses.
From the combined fit, a neutrino signal of () events in the reactor on data set was obtained. The significance of this event excess corresponds to . The neutrino signal at low energies of the spectrum is illustrated in Fig.3. Fits using just single detectors independently give consistent results. The result was cross-checked in two implementations of the likelihood fit and agreed within less than 2%. Additional systematic uncertainties related to background model, non-linearity correction and fit systematics were studied independently and are included in the fit result.
@@
The Conus+ signal of () measured neutrinos is fully consistent with the expectation of () events. This implies agreement of the Conus+ data with the CENS cross-section of the SM and the estimated antineutrino flux based on the thermal power of the reactor. Moreover, the detected rate is in very good agreement with the predicted Ge quenching using the Lindhard theory with a quenching parameter as measured in [34]. The deviations from Lindhard theory claimed in [33] and the excess reported in [17], which is based on it, are both ruled out by this result. The claim was already disfavored before by the independent measurement of the quenching factor [34], previous Conus results [27] and theoretical considerations [36].
@@
In Table 1, the ratio of measured and predicted neutrino interactions is shown for the Conus+ result. This ratio is compared to the ratios measured in the COHERENT measurements at higher neutrino energies using multiple target nuclei and the results from coherent scattering of solar neutrinos. The use of different target nuclei allows to study the quadratic enhancement of the cross-section by the number of neutrons in the target. In the COHERENT data for Ge, the measured rate was found slightly below the predicted value, although not highly significant. Such a deficit was not confirmed in the Conus+ data. The combination of the CENS result of Conus+ with results of the germanium target data of the COHERENT experiment can in principle be used to extract information about nuclear form factors in neutrino light.
After the first detection of CENS at a nuclear reactor, as reported in this article, the next step will be a more precise measurement of the CENS cross-section. Higher precision can be achieved by increasing the target mass, lowering the energy threshold of the detectors and longer operation times, in particular in the reactor off phases. Therefore, 3 of the 4 Conus+ detectors were replaced in November 2024 by a newer generation. These new detectors have a larger mass of 2.4 kg each. First characterizations in the MPIK laboratory indicated even lower energy thresholds. The detector with the lowest energy threshold in the first Conus+ run (C3) was kept as a reference for a better comparison between the phases of the experiment. With this detector configuration it is planned to measure for another few years.
A high-statistics CENS measurement might open a new phase in fundamental physics and will allow to study physics within and beyond the SM. The measured CENS rate can for example be affected by new mediator particles similar to the Z boson, electromagnetic properties of neutrinos, or non-standard interactions. Moreover, it is possible to study the Weinberg angle at low energies, the existence of sterile neutrinos or supernovae astrophysics. There are also connections to dark matter experiments. A precise CENS measurement will allow to learn more about the neutrino sources as the Sun or nuclear reactors. The evolution of the reactor thermal power and fissile isotope concentrations in fuel elements could be monitored with rather small and mobile neutrino detectors. In summary, there is a wide range of topics ranging from BSM theories, nuclear physics and astrophysics that can be addressed with CENS measurements at nuclear reactors. The Conus and Conus+ experiments are pioneering in this field.